### v1

@@ 4 Results
In order to assess the performance of our modelling framework, we considered three alternative model specifications alongside the main model presented in Section 3. These variants differ in two aspects: (i) the treatment of the gender effect, modelled either as a fixed or a random effect; and (ii) the prior distribution on the auto-regressive coefficient $\rho$, specified either as a Beta(1,1)(1,1) distribution (constraining $\rho$ to the unit interval) or as a standard Normal distribution (allowing unrestricted support). This results in four model configurations, named M1, M2, M3 and M4, as detailed in Table 11 in Appendix 12. Note that M1 (i.e., gender as fixed effect and Beta prior for $\rho$) corresponds to our main model described in Section 3. To compare the predictive performance of these four models, we performed a temporal training-test split, using data from 2022 ($t=13$) as the test set and the previous years ($t=1,\dots,12$) for training. Models were assessed using standard predictive metrics. The results indicate that M1 outperforms the competing models in terms of predictive accuracy (see Table 12 in Appendix 12). The latter is likely due to the temporal dynamics allowed by the Normal prior, which can lead to explosive behaviours and overfitting. The random gender effect offered no improvement over the fixed-effect specifications. Overall, M1 offered the best balance between parsimony and predictive accuracy.
5 Discussion
@@
The choice of the Beta distribution is motivated not only by its support on the unit interval - making it particularly suitable for modelling rates and proportions - but also by its flexibility, as it accommodates a wide range of distributional shapes, depending on the parameter values (Johnson et al.,, 1995). In the present case study, the proposed Bayesian hierarchical Beta regression model incorporates structured spatial and temporal random effects additively, along with gender-specific intercepts. This model proves effective in capturing regional variations and temporal dependencies in obesity rates.
@@
Some methodological limitations warrant consideration. The ICAR prior for the structured spatial random effects assumes complete spatial dependence, which may not fully align with the underlying spatial processes. Additionally, the additive specification of spatial and temporal random effects with separate priors may overlook potential space-time interactions. The exclusion of unstructured spatio-temporal variability, which could enhance flexibility in capturing localized deviations, represents another limitation. Furthermore, the regional-level analysis improves measurement reliability and mitigates data gaps but may obscure finer-scale variations at the provincial level. Despite these constraints, the study tries to advance the robustness and interpretability of spatio-temporal models, prioritizing methodological clarity and practical applicability. It contributes to the development of statistical tools for spatio-temporal analysis in public health and social sciences, offering a comprehensive understanding of obesity trends across Italian regions.
@@
Our findings highlight the substantial role of the gender-specific random effects in estimating obesity outcomes, underscoring gender as a critical determinant of regional obesity rates. Spatial random effects reveal homogeneous patterns across the North-West, with the exception of Emilia-Romagna, which exhibits higher values. In the North-East, Friuli-Venezia Giulia and Veneto show similar values, while Trentino-Alto Adige displays lower magnitudes with a greater range. The Centre geographical zone is characterized by marked heterogeneity, with Lazio showing significantly lower values and Umbria higher ones. Southern regions generally exhibit higher spatial random effects, except for Calabria. The Islands show distinct spatial structures, an aspect enforced by the fact that Sicilia is treated as a neighbour to Calabria in the ICAR prior, while Sardegna is considered spatially isolated. Temporal random effects indicate a stable trend in the early years, followed by an increase in later years, suggesting a growing upward trend in obesity rates. Fixed effects, represented by exogenous covariates, appear to play a minimal role. Indeed, the application of the SSVS prior identifies only few covariates with significant associations to obesity, namely life expectancy, life expectancy in good health and overweight.
While this study does not establish causal relationships, it underscores the importance of modelling obesity trends through demographic and spatio-temporal dynamics rather than relying solely on fixed socioeconomic or lifestyle predictors. The findings reinforce the necessity of incorporating spatial and temporal structures in public health research. Future research could refine this approach by analysing finer spatial resolutions, such as provinces or municipalities, where fixed effects may be more relevant. Incorporating individual-level data and environmental factors could further enhance understanding of obesity dynamics, while non-parametric methods could facilitate data-driven regional clustering. Additionally, exploring interactions between spatial, temporal, and other random effects could provide deeper insights into regional health patterns.
In conclusion, this study demonstrates the effectiveness of spatio-temporal Bayesian modelling in capturing the complexities of obesity rates and in informing societal challenges using administrative data. It advocates for continued investigation of demographic, geographic, temporal, and behavioural factors to inform public health interventions and policy-making.
6 Competing interests
The authors declare that they have no competing interests.
7 Author contributions statement
L.R.: data collection and engineering, methodology, formal analysis, writing - original draft. R.A.: methodology, supervision, review & editing. M.C.: methodology, supervision, review & editing.
8 Data and code
The assembled dataset and the code used for implementing the analysis and running the models is available at the following GitHub repository:
https://github.com/LucianoRota00/Spatio-Temporal-Dynamics-of-Obesity-in-Italian-Regions.git.
9 Acknowledgments
L.R. gratefully acknowledges the financial support provided by the University of Bergamo for the Ph.D. programme in Statistics at the University of Milano-Bicocca.
{appendices}
10 Appendix A: description of the covariates
@@
Table 7: Description of the covariates used in the analysis, grouped into four thematic categories: Socio-Demography, Mortality, Lifestyle, and General Wealth-related Information. The table reports the unit of measurement, availability by gender, and data source for each variable.
@@
Health expenditure per capita euro/population no MEF
@@
{tablenotes}
Source: HFA = Health for All database; BES = Italian Equitable and Sustainable Well-being report; MEF = Italian Ministry of Economy and Finance; Eurostat = Statistical Office of the European Union.
11 Appendix B: posterior summary statistics
This appendix provides the posterior summary statistics not reported in the main text.
Table 8: Posterior summary statistics for the region-specific spatial parameters $\psi_{i}$, with $i=1,\ldots,20$.
. Parameter (Region) Mean SD CV Skewness Kurtosis $\psi_{1}$ (Piemonte) -0.61 0.07 -0.12 0.18 3.46 $\psi_{2}$ (Valle d’Aosta) -0.48 0.09 -0.18 0.34 3.96 $\psi_{3}$ (Lombardia) -0.41 0.08 -0.19 0.25 3.41 $\psi_{4}$ (Liguria) -0.62 0.08 -0.12 0.11 3.42 $\psi_{5}$ (Emilia-Romagna) 0.25 0.08 0.33 0.32 3.59 $\psi_{6}$ (Trentino-Alto Adige) -0.49 0.12 -0.25 0.02 3.19 $\psi_{7}$ (Veneto) -0.02 0.07 -3.34 0.06 3.13 $\psi_{8}$ (Friuli-Venezia Giulia) 0.13 0.07 0.55 0.03 3.08 $\psi_{9}$ (Toscana) -0.15 0.08 -0.56 0.18 3.30 $\psi_{10}$ (Umbria) 0.29 0.07 0.26 0.19 3.29 $\psi_{11}$ (Marche) 0.03 0.07 2.73 0.03 3.09 $\psi_{12}$ (Lazio) -0.40 0.07 -0.18 0.20 3.55 $\psi_{13}$ (Abruzzo) 0.49 0.06 0.13 -0.03 3.05 $\psi_{14}$ (Molise) 0.89 0.08 0.09 -0.25 3.57 $\psi_{15}$ (Campania) 0.35 0.11 0.32 -0.31 3.65 $\psi_{16}$ (Puglia) 0.65 0.08 0.12 -0.55 4.34 $\psi_{17}$ (Basilicata) 0.70 0.09 0.13 -0.26 3.58 $\psi_{18}$ (Calabria) -0.07 0.12 -1.69 -0.13 3.12 $\psi_{19}$ (Sicilia) 0.16 0.10 0.60 -0.22 3.60 $\psi_{20}$ (Sardegna) -0.71 0.09 -0.12 -0.32 3.79
Table 9: Posterior summary statistics for time-specific parameters $\alpha_{t}$, with $t=1,\ldots,13$.
Parameter (Year) Mean SD CV Skewness Kurtosis
@@ 11 Appendix B: posterior summary statistics
Table 10: Posterior summary statistics for precisions of the spatial and temporal random effects, $\tau_{\psi}$ and $\tau_{\alpha}$, respectively.
@@
[…]
Focusing on the precision parameters for the spatial and temporal random effects ($\tau_{\psi}$ and $\tau_{\alpha}$, respectively), we observe substantial Bayesian updating, particularly for the temporal component. Beginning with conservative Gamma priors, the posterior distribution of the spatial precision $\tau_{\psi}$ exhibits only a moderate increase. This restrained growth can be partly attributed to the ICAR prior structure, in which the full-conditional precision is scaled by the number of neighboring regions. In contrast, the temporal precision $\tau_{\alpha}$, whose prior specification does not involve such a multiplicative adjustment, undergoes a more pronounced shift, with a posterior mean markedly higher than that of $\tau_{\psi}$. However, direct comparison between these two parameters should be approached with caution, as their prior formulations differ substantially, rendering such comparisons potentially misleading. Both posterior distributions are characterized by high skewness and kurtosis, suggesting strong positive asymmetry and heavy tails.
12 Appendix C: predictive performance comparison
@@
where $\beta_{0}$ is the global intercept and $\gamma_{s}$ represents a random deviation for each gender. The random effects follow a classical hierarchical formulation (e.g., Hoff,, 2009; Gelman,, 2006):
@@
In contrast, the gende fixed effect specification is the one we describe in Section 3.2.
As for the auto-regressive coefficient $\rho$, we consider two prior formulations: a $\text{Beta}(1,1)$ prior (constraining $\rho$ to the interval $(0,1)$) as specified in Section 3.2, and a standard normal prior $\mathcal{N}(0,1)$, which places no support restrictions and permits values outside the unit interval.
@@
Table 12: Predictive performance metrics for the four model variants, evaluated on the 2022 test data. RMSE (F) and RMSE (M) refer to female and male subgroups, respectively.

### v2

@@ 4.1 Predictive performance
In order to assess the performance of the modelling approach we propose, we considered three alternative model specifications alongside the main model presented in Section 3. These variants differ in two aspects: (i) the treatment of the gender effect, modelled either as a fixed or a random effect; and (ii) the prior distribution on the auto-regressive coefficient $\rho$, specified either as a Beta(1,1)(1,1) distribution (constraining $\rho$ to the unit interval) or as a standard Normal distribution (allowing unrestricted support). This results in four model configurations, named M1, M2, M3 and M4, as detailed in Table 11 in Appendix D.1. Note that M1 (i.e., gender as fixed effect and Beta prior for $\rho$) corresponds to our main model described in Section 3. To evaluate and compare the predictive performance of the four models, we implemented both temporal and spatial validation strategies. First, we conducted a temporal train–test split, using data from 2022 ($t=13$) as the test set and data from the preceding years ($t=1,\dots,12$) for model training (see Appendix D.1). Second, we performed a spatial cross-validation procedure by leaving out one representative region from each macro-area: Lombardia (North), Toscana (Centre), and Campania (South). Further details are provided in Appendix D.2. Models were assessed using standard predictive metrics. For what concerns the temporal test, the results reported in Table 12 indicate that M1 outperforms the competing models in terms of predictive accuracy. The latter is likely due to the temporal dynamics allowed by the Normal prior, which can lead to explosive behaviours and overfitting. The random gender effect offered no improvement over the fixed-effect specifications. Overall, M1 offered the best balance between parsimony and predictive accuracy. For what concerns the spatial cross-validation, we obtained smaller performance differences across model specifications than those observed in the temporal validation exercise. This indicates that, once trained on a sufficiently broad set of regions, all four models are able to capture the main spatial patterns in obesity prevalence with comparable accuracy.
5 Concluding remarks
@@
The choice of the Beta distribution is motivated not only by its support on the unit interval - making it particularly suitable for modelling rates and proportions - but also by its flexibility, as it accommodates a wide range of distributional shapes, depending on the parameter values [23]. We mention that beyond Beta regression, alternative methodologies for analysing rate or proportion data have been proposed. For example, compositional data analysis has been used to jointly model multiple outcome categories, such as underweight, overweight, and obesity, through log-ratio transformations [28]. While such approaches are valuable when the primary interest lies in the joint dynamics of multiple components, our focus is on modelling a single rate outcome over space and time. In the present case study, the proposed Bayesian hierarchical Beta regression model incorporates structured spatial and temporal random effects additively, along with gender-specific intercepts. This model proves effective in capturing regional variations and temporal dependencies in obesity rates. While the individual components of the proposed model are well established, their integration into a unified Bayesian hierarchical framework constitutes the main methodological contribution of this work, offering a coherent and flexible approach for analysing regional obesity rates and their spatio-temporal dynamics.
@@
Some methodological limitations warrant consideration. The ICAR prior for the structured spatial random effects assumes complete spatial dependence, which may not fully reflect the underlying spatial processes. Additionally, the additive specification of spatial and temporal random effects, with separate priors, may overlook potential space-time interactions. The exclusion of unstructured spatio-temporal variability, which could increase flexibility in capturing localized deviations, represents a further limitation. Furthermore, although the regional-level analysis improves measurement reliability and mitigates data gaps, it may obscure finer-scale variations at the provincial level. Finally, the adjacency matrix assigns equal weight only to immediately neighbouring regions and therefore does not account for distance decay or other forms of connectivity, such as economic or cultural links. While this is a standard and interpretable choice for areal data, future work could explore more flexible weighting schemes to better capture complex spatial relationships. Despite these limitations, the proposed framework remains parsimonious, interpretable, and effective in predictive terms. Its main strength lies in balancing modelling flexibility with methodological clarity, thereby providing a practically useful approach for the spatio-temporal analysis of regional obesity rates in public health applications.
@@
Our findings highlight the substantial role of the gender-specific random effects in estimating obesity outcomes, underscoring gender as a critical determinant of regional obesity rates. Spatial random effects reveal homogeneous patterns across the North-West, with the exception of Emilia-Romagna, which exhibits higher values. In the North-East, Friuli-Venezia Giulia and Veneto show similar values, while Trentino-Alto Adige displays lower magnitudes with a greater range. The Centre geographical zone is characterized by marked heterogeneity, with Lazio showing significantly lower values and Umbria showing higher ones. Southern regions generally exhibit higher spatial random effects, except for Calabria. The Islands show distinct spatial structures, an aspect enforced by the fact that Sicilia is treated as a neighbour to Calabria in the ICAR prior, while Sardegna is considered spatially isolated. Temporal random effects indicate a stable trend in the early years, followed by an increase in later years, suggesting a growing upward trend in obesity rates. Under the SSVS prior specification, most coefficients are shrunk toward zero, with only a few covariates being identified as having a relevant association with obesity, namely “life expectancy”, “life expectancy in good health”, and “overweight”. This result suggests that spatial and temporal random effects explain a substantial portion of the variability in obesity rates. Clearly, additional factors, such as environmental characteristics, food environment indicators, or policy-related variables, may play a relevant role but could not be considered due to the limited availability of regional-level data over the study period. The integration of richer covariate information, therefore, represents a natural and promising direction for future research as more comprehensive data become accessible.
While this study does not aim to establish causal relationships, it highlights the importance of modelling obesity trends through demographic and spatio-temporal structures rather than relying solely on fixed socioeconomic or lifestyle predictors. The results further support the value of incorporating spatio-temporal dependence in public health research.
Future research could extend this framework by considering finer spatial resolutions, such as provinces or municipalities, incorporating individual-level or environmental information, and allowing for spatio-temporal interactions. Alternative prior specifications could also be explored to increase flexibility in the presence of outliers or asymmetric variation. More broadly, the proposed Bayesian hierarchical Beta regression model may be useful for other areal outcomes bounded between 0 and 1 and characterized by spatial and temporal dependence, including applications in public health and the social sciences.
Acknowledgements
Blinded for review
Author contribution
Blinded for review
Funding
Not applicable
Competing interests
Not applicable
Ethics approval and consent to participate
Not applicable
Consent for publication
Not applicable
Appendix A Appendix Organization
We provide additional material in the appendices: Appendix B contains a table reporting the classification of Italy into geographical zones and a table describing the covariates used in the analysis; Appendix C reports additional posterior summary statistics not included in the main text for brevity, Appendix D presents the results of the predictive comparison of some variants of the main model, Appendix E includes sensitivity analyses for different SSVS hyperparameter specifications and, finally, Appendix F provides extensive diagnostics and performance metrics.
Appendix B Geographical Areas and Covariates
Zone Regions
North East (NE) Trentino-Alto Adige, Veneto, Friuli-Venezia Giulia
Centre (C) Toscana, Umbria, Marche, Lazio
South (S) Abruzzo, Molise, Campania, Puglia, Basilicata, Calabria
Islands (SI) Sicilia, Sardegna
[…]
@@ Appendix B Geographical Areas and Covariates
Table 7: Covariates used in the analysis, grouped into four thematic categories: Socio-Demography, Mortality, Lifestyle, and General Wealth-related Information. The table reports the unit of measurement, availability by gender, and data source for each variable.77footnotetext: Source: HFA = Health for All database; BES = Italian Equitable and Sustainable Well-being report; MEF = Italian Ministry of Economy and Finance; Eurostat = Statistical Office of the European Union.
@@
Health expenditure per capita euro/population yes MEF
@@
Appendix C Posterior Summary Statistics
This appendix provides the parameter posterior summary statistics not reported in the main text.
Table 8: Posterior summary statistics for region-specific parameters $\psi_{1}$ to $\psi_{20}$.
Parameter (Region) Mean SD CV Skewness Kurtosis
$\psi_{1}$ (Piemonte) -0.61 0.07 -0.12 0.18 3.46
$\psi_{2}$ (Valle d’Aosta) -0.48 0.09 -0.18 0.34 3.96
$\psi_{3}$ (Lombardia) -0.41 0.08 -0.19 0.25 3.41
$\psi_{4}$ (Liguria) -0.62 0.08 -0.12 0.11 3.42
$\psi_{5}$ (Emilia-Romagna) 0.25 0.08 0.33 0.32 3.59
$\psi_{6}$ (Trentino-Alto Adige) -0.49 0.12 -0.25 0.02 3.19
$\psi_{7}$ (Veneto) -0.02 0.07 -3.34 0.06 3.13
$\psi_{8}$ (Friuli-Venezia Giulia) 0.13 0.07 0.55 0.03 3.08
$\psi_{9}$ (Toscana) -0.15 0.08 -0.56 0.18 3.30
$\psi_{10}$ (Umbria) 0.29 0.07 0.26 0.19 3.29
$\psi_{11}$ (Marche) 0.03 0.07 2.73 0.03 3.09
$\psi_{12}$ (Lazio) -0.40 0.07 -0.18 0.20 3.55
$\psi_{13}$ (Abruzzo) 0.49 0.06 0.13 -0.03 3.05
$\psi_{14}$ (Molise) 0.89 0.08 0.09 -0.25 3.57
$\psi_{15}$ (Campania) 0.35 0.11 0.32 -0.31 3.65
$\psi_{16}$ (Puglia) 0.65 0.08 0.12 -0.55 4.34
$\psi_{17}$ (Basilicata) 0.70 0.09 0.13 -0.26 3.58
$\psi_{18}$ (Calabria) -0.07 0.12 -1.69 -0.13 3.12
$\psi_{19}$ (Sicilia) 0.16 0.10 0.60 -0.22 3.60
$\psi_{20}$ (Sardegna) -0.71 0.09 -0.12 -0.32 3.79
Table 9: Posterior summary statistics for time-specific parameters $\alpha_{1}$ to $\alpha_{13}$.
Parameter (Year Mean SD CV Skewness Kurtosis
@@ Appendix C Posterior Summary Statistics
Table 10: Posterior summary statistics for the hyperparameters of space ($\tau_{\psi}$), and time ($\tau_{\alpha}$) random effects
@@
Focusing on the precision parameters for the spatial and temporal random effects ($\tau_{\psi}$ and $\tau_{\alpha}$, respectively), we observe substantial Bayesian updating, particularly for the temporal component (see Table 10). Beginning with conservative Gamma priors, the posterior distribution of the spatial precision $\tau_{\psi}$ exhibits only a moderate increase. This restrained growth can be partly attributed to the ICAR prior structure, in which the full-conditional precision is scaled by the number of neighboring regions. In contrast, the temporal precision $\tau_{\alpha}$, whose prior specification does not involve such a multiplicative adjustment, undergoes a more pronounced shift, with a posterior mean markedly higher than that of $\tau_{\psi}$. However, direct comparison between these two parameters should be approached with caution, as their prior formulations differ substantially, rendering such comparisons potentially misleading. Both posterior distributions are characterized by high skewness and kurtosis, suggesting strong positive asymmetry and heavy tails.
Appendix D Predictive Performance Comparison
D.1 Predictive Test for Year 2022
@@
where $\beta_{0}$ is the global intercept and $\gamma_{s}$ represents a random deviation for each gender. The random effects follow a classical hierarchical formulation [e.g., 21, 15]:
@@
In contrast, the gender fixed effect specification is the one we describe in Section 3.2.
As for the autoregressive coefficient $\rho$, we consider two prior formulations: a $\text{Beta}(1,1)$ prior (constraining $\rho$ to the interval $(0,1)$) as specified in Section 3.2, and a standard normal prior $\mathcal{N}(0,1)$, which places no support restrictions and permits values outside the unit interval.
@@
Table 12: Predictive performance metrics for the four model variants, evaluated on 2022 test data. RMSE (F) and RMSE (M) refer to female and male subgroups, respectively.